I am in year 7 and I have a question. Could you please explain the term Linear Equation. I have looked on the internet, but the explanations are too complicated, could you please explain it to me so it can make more sense. Thank You, Ryan aged 12
Answer by AnlytcPhil(1116)

I am in year 7 and I have a question. Could you please explain the term
Linear Equation. I have looked on the internet, but the explanations are
too complicated, could you please explain it to me so it can make more
sense. Thank You, Ryan aged 12
`
The first 4 letters of the word "LINEar" spell "LINE". So an equation in
y and x which has a graph which is a straight line, and not a curve, is
called a "linear equation". It turns out that a linear equation is one
that when it is simplified, has no x² or x³ or higher powers of x than
the first power, no denominators containing x, no square roots containing
x, no cube roots containing x, etc. Here are some examples of linear
equations:
`
5x - 4y = 9
`
3y = 4x + 2
`
y = -7x - 2
`
(3/4)x + (1/3)y = 1/2
`
Any equation that can be put in the form
`
Ax + By = C
`
where A, B and C are numbers is a linear equation, i.e., its graph is
a straight line.
`
Edwin J